# On BRST Lagrangian description

### v1

@@
We present an exhaustive BRST lagrangian description of partially massless bosonic fields in four-dimensional space. The basic fields are formulated in terms of two-component spin-tensors in (A)dS space where the tracelessness conditions are automatically fulfilled. The mass shell of partially massless fields is reformulated in terms of constraints on Fock space vectors including the second-class constraints. A conversion procedure for transforming second-class constraints into first-class ones is developed, allowing one to construct a Hermitian and nilpotent BRST charge in the Fock space under consideration. It is proven that the hermiticity and nilpotency restrict the conditions on the theory parameters, which are fulfilled only in dS space. The hermiticity of the BRST charge is incompatible with AdS space. The gauge invariant Lagrangian is constructed on the basis of the BRST charge, and for spin $s$ and depth $t$ the allowed states in the Lagrangian include only (st)(s-t) Stückelberg fields. Their exclusion leads to gauge transformations of degree (st)(s-t) for the physical fields. The Lagrangian equations of motion exactly reproduce the mass shell conditions. The Lagrangian in terms of conventional spin-tensor fields is also presented.
@@ 1 Introduction
The discovery of partially massless fields sparked a surge of interest in constructing appropriate representations of the (A)dS group in different dimensions and in developing methods for describing these fields including Lagrangian formulation, supersymmetric issues and various applications [17, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. We emphasize that partially massless fields are defined only in (A)dS spaces, and the aspects of unitarity of the corresponding field theory are highly nontrivial compared to flat space theories and need to be specially clarified.
@@ 4.2 Construction of the converted constraints and the Hermitian and nilpotent BRST charge
To conclude this section, let us emphasize some properties of the operator $A(\mathrm{n})$ at $m^{2}=m^{2}_{t}$ (4.40) in dS space. It is easy to see that the vectors $|\phi_{k}\rangle$ (4.42) are the eigenvectors of this operator with eigenvalues
@@ 5.2 Derivation of mass-shell conditions
Gauge transformations of the other vectors $|\Phi_{s}\rangle$ and $|{\mathcal{X}}_{s-2}\rangle$ (5.10)) with the confined parameter $|\Lambda_{s-1}\rangle$ for the component vectors $\varphi_{s}$ and $\varphi_{k}\rangle$ for t<k<s take the form
@@
$\delta|\varphi_{k}\rangle$ $=$ $l^{+}|\lambda_{k-1}\rangle+A(s{-}k{-}1)\;|\lambda_{k}\rangle\,,\qquad t<k<s\,.$ (5.21)
@@
on the gauge parameters. After that, using gauge transformations $\delta|\varphi_{t+2}\rangle=l^{+}|\lambda_{t+1}\rangle+A(s-t-1)\;|\lambda_{t+2}\rangle$ from (5.21), we can impose the gauge fixing condition $|\varphi_{t+2}\rangle=0$. The residual gauge transformations that preserve this gauge fixing condition must be subject to the condition
$|\lambda_{t+2}\rangle={}-\Bigl[A(s-t-3)\Bigr]^{-1}l^{+}|\lambda_{t+1}\rangle=\Bigl[A(s-t-2)A(s-t-3)\Bigr]^{-1}l^{+}|\lambda_{t}\rangle$ (5.23)
@@ 6 Component Lagrangian
$\langle 0|a_{\alpha_{1}}\ldots a_{\alpha_{k}}\bar{a}^{\dot{\alpha}_{1}}\ldots\bar{a}^{\dot{\alpha}_{k}}b^{\beta_{1}}\ldots b^{\beta_{n}}\bar{b}_{\dot{\beta}_{1}}\ldots\bar{b}_{\dot{\beta}_{n}}|0\rangle={}\delta_{kn}(k!)^{2}\delta_{(\alpha_{1}}^{\beta_{1}}\ldots\delta_{\alpha_{k})}^{\beta_{k}}\delta^{(\dot{\alpha}_{1}}_{\dot{\beta}_{1}}\ldots\delta^{\dot{\alpha}_{k})}_{\dot{\beta}_{k}}\,$. (6.2)
@@ 7 Summary
  • The hermitianity of the BRST charge in the theory under consideration requires that the vectors used be restricted to expansions (4.46) and (4.47). Moreover, the presence of (st)(s-t) Stückelberg fields in expansion (4.46) leads, after their exclusion, to gauge transformations (5.27) of physical fields of the (st)(s-t) order in derivatives.
@@ Acknowledgments
The authors are grateful to Yurii M. Zinoviev for discussion and useful comments.

### v3

@@
Dedicated to the bright memory of I.V. Tyutin
We present an exhaustive BRST lagrangian description of partially massless bosonic fields in four-dimensional space. The basic fields are formulated in terms of two-component spin-tensors in (A)dS space where the tracelessness conditions are automatically fulfilled. The mass shell of partially massless fields is reformulated in terms of constraints on Fock space vectors including the second-class constraints. A conversion procedure for transforming second-class constraints into first-class ones is developed, allowing one to construct a Hermitian and nilpotent BRST charge in the Fock space under consideration. It is proven that the hermiticity and nilpotency restrict the conditions on the theory parameters, which are fulfilled only in dS space. The hermiticity of theBRST charge is incompatible with AdS space. The gauge invariant Lagrangian is constructed on the basis of the BRST charge, and for spin $s$ and depth $t$ the allowed states in the Lagrangian include only (st1)(s-t-1) Stückelberg fields. Their exclusion leads to gauge transformations of degree (st)(s-t) for the physical fields. The Lagrangian equations of motion exactly reproduce the mass shell conditions. The Lagrangian in terms of conventional spin-tensor fields is also presented.
@@ 1 Introduction
The discovery of partially massless fields sparked a surge of interest in constructing appropriate representations of the (A)dS group in different dimensions and in developing methods for describing these fields including Lagrangian formulation, supersymmetric issues and various applications [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. We emphasize that partially massless fields are defined only in (A)dS spaces, and the aspects of unitarity of the corresponding field theory are highly nontrivial compared to flat space theories and need to be specially clarified.
@@ 4.2 Construction of the converted constraints and the Hermitian and nilpotent BRST charge
To conclude this section, let us emphasize some properties of the operator $A(\mathrm{n})$ at $m^{2}=m^{2}_{t}$ (4.40) in dS space. It is easy to see that the vectors $(b^{+})^{k}|\varphi_{l}\rangle$ at any $l$ are the eigenvectors of this operator with eigenvalues
@@ 5.2 Derivation of mass-shell conditions
Gauge transformations of the other vectors $|\Phi_{s}\rangle$ and $|{\mathcal{X}}_{s-2}\rangle$ (5.10)) with the confined parameter $|\Lambda_{s-1}\rangle$ for the component vectors $|\varphi_{s}\rangle$ and $|\varphi_{k}\rangle$ for t<k<s take the form
@@
$\delta|\varphi_{k}\rangle$ $=$ $l^{+}|\lambda_{k-1}\rangle+A(s{-}k{-}1)\;|\lambda_{k}\rangle\,,\qquad t<k<s\,$, (5.21)
@@
on the gauge parameters. After that, using gauge transformations $\delta|\varphi_{t+2}\rangle=l^{+}|\lambda_{t+1}\rangle+A(s-t-3)\;|\lambda_{t+2}\rangle$ from (5.21), we can impose the gauge fixing condition $|\varphi_{t+2}\rangle=0$. The residual gauge transformations that preserve this gauge fixing condition must be subject to the condition
$|\lambda_{t+2}\rangle={}-\Bigl[A(s-t-3)\Bigr]^{-1}l^{+}|\lambda_{t+1}\rangle=\Bigl[A(s-t-2)A(s-t-3)\Bigr]^{-1}(l^{+})^{2}|\lambda_{t}\rangle$ (5.23)
@@ 6 Component Lagrangian
$\langle 0|a_{\alpha_{1}}\ldots a_{\alpha_{k}}\bar{a}^{\dot{\alpha}_{1}}\ldots\bar{a}^{\dot{\alpha}_{k}}c^{\beta_{1}}\ldots c^{\beta_{n}}\bar{c}_{\dot{\beta}_{1}}\ldots\bar{c}_{\dot{\beta}_{n}}|0\rangle={}\delta_{kn}(k!)^{2}\delta_{(\alpha_{1}}^{\beta_{1}}\ldots\delta_{\alpha_{k})}^{\beta_{k}}\delta^{(\dot{\alpha}_{1}}_{\dot{\beta}_{1}}\ldots\delta^{\dot{\alpha}_{k})}_{\dot{\beta}_{k}}\,$. (6.2)
@@ 7 Summary
  • The hermitianity of the BRST charge in the theory under consideration requires that the vectors used be restricted to expansions (4.46) and (4.47). Moreover, the presence of (st1)(s-t-1) Stückelberg fields in expansion (4.46) leads, after their exclusion, to gauge transformations (5.27) of physical fields of the (st)(s-t) order in derivatives.
@@ Acknowledgments
The authors are grateful to K.B. Alkalaev, A.O. Barvinsky, M.A. Vasiliev, B.L. Voronov and Yu.M. Zinoviev for discussion and useful comments.